OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a_1 = 15 and a_13= -57?

OpenStudy (anonymous):

here u need to first find the common difference of the arithmetic sequence using a_n=a_1 +(n-1)*d thus we have -57 = 15 +(13-1)*d or d= -72/12 =-6 hence a_32=15 +(32-1)*(-6)=15-186=-171